OpenStudy (anonymous):

PLease Please Please help A line segment, ST, has endpoints S(-4,-5) and T(2,-3). Which of the following equations represents the perpendicular bisector of the line segment?

OpenStudy (anonymous):

y = -3x - 7 y = 2x - 8 y = 6x - 7 y = -3x - 13

OpenStudy (anonymous):

I got this one sec

OpenStudy (anonymous):

y = -3x - 7 is the answer.

OpenStudy (anonymous):

thank you can you help me with another

OpenStudy (anonymous):

Need an explanation?

OpenStudy (anonymous):

sure if you medal me bud ;)

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

What's the next question

OpenStudy (anonymous):

The answer is A.

OpenStudy (anonymous):

thanks..... can you do another

OpenStudy (anonymous):

For these problems, all you do is find the midpoint of the two coordinates then find the oppositte recriprical of the slope, and solve

OpenStudy (anonymous):

Sure I guess

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

My bad it's C

OpenStudy (anonymous):

Oppositte recriprocal slopes, therefore are perpendicular

OpenStudy (anonymous):

thank you so so so much :)

OpenStudy (anonymous):

No problem bud, keep practicing